# Safety Filters for Black-Box Dynamical Systems by Learning Discriminating Hyperplanes

### v1

@@ 4.1 Supervised learning approach
For $\Delta t>0$, there exists $a:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and $b:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that are Lipschitz continuous in $x$, such that $\Pi(x)=\{u\in U\;|\;a(x)^{\top}u\geq b(x)\}\subseteq\{u\in U\;|\;I_{S}(\mathrm{x}(\Delta t))=1\}$ where x()normal-$\mathrm{x}(\cdot)$ is from (1) with $\mathrm{x}(0)=x,\mathrm{u}(\cdot)\equiv u$. Moreover, for small enough $\Delta t$, $\Pi(x)$ is not empty if $\{u\in U\;|\;I_{S}(\mathrm{x}(\Delta t))=1\}$ has a non-empty interior. (Proof: Appendix 7.2)
By using the lookahead approach, $\Pi(x)$ satisfying Theorem 2 anticipates into the future, which is also the main mechanism of CBFs (Choi et al.(2023)Choi, Lee, Li, How, Sreenath, Herbert, and Tomlin) or predictive filter (Wabersich and Zeilinger(2021)). If the lookahead time is small, then the constraint would be active only on states that are close to the boundary. In contrast, larger $\Delta t$ will encode additional safety margin from the boundary of $S$, however, it might make $\Pi(x)$ an empty set. In Section 5, we demonstrate how this design results in a “smooth braking” behavior as $\mathrm{x}(\cdot)$ approaches the boundary of the safe domain, similarly to the CBF-based design, and how different choice of the lookahead time affects this behavior.
Next, we use this idea to learn the discriminating hyperplane with a neural network. We train $[a_{\theta}(x),b_{\theta}(x)]\in\mathbb{R}^{m+1}$, where $\theta$ is the neural network weights, in a supervised manner using a dataset of state and input transition pairs, and their labels detailing if an input is safe. Specifically, $N$ states, $\{x_{i}\}_{i=1}^{N}$, are uniformly sampled from the control invariant set $S$. Then $M$ inputs per state, sampled uniformly from $U$, $\{u_{ij}\}_{j=1}^{M}$, are applied to the dynamical system by sample-and-hold for the lookahead time $\Delta t$, resulting in the terminal state $\mathrm{x}_{ij}(\Delta t)$. The indicator function of the set $S$ at the terminal state, $I_{S}(\mathrm{x}_{ij}(\Delta t))$, determines the label of whether the next state is safe or unsafe:
@@
If the prediction of safety from the neural network matches the labels, the loss is not incurred. False positive samples whose unsafe control input is predicted to be safe are counted in the first term, and false negative samples whose safe control input is predicted to be unsafe are counted in the second term. It is worth noting that the structure of the loss exhibits similarities with the loss design employed in Dawson et al.(2023)Dawson, Gao, and Fan for learning certificate functions.
@@ 4.2 Reinforcement learning approach
An alternative approach is to employ RL, when it is hard to find a control invariant set $S$ for the supervised learning method. We propose a method analogous to PPO, where instead of learning a policy that maximizes reward, we use an actor parametrizing the discriminating hyperplane to minimize instances of safety violation. Our method uses the form $\pi^{a}_{\theta}(a|x),\pi^{b}_{\theta}(b|x)$ as two normal distributions to sample the hyperplane parameters. During the rollout, based on the sampled $(a,b)$ for each $x$, with probability $1\!-\!\delta$, we sample $u$ from $\{u\in U|\;au\geq b\}$ which enforces the learned safety constraint, and with probability $\delta$, we sample $u$ from $U$, allowing an exploration with a small probability. Given the target constraint set $X$, the reward for learning the hyperplane is designed as
$r(x,u)=\begin{cases}1+d(u)&\text{if }x\in X,\\ c&\text{otherwise,}\end{cases}$ (7)
where $c\leq 0$ induces a negative reward when safety is violated and $d(u)$ is an optional bonus term that can add to the reward if the input $u$ is in the action space. The reward is positive when safety is satisfied and the bonus term incentivizes the hyperplane to only constrain the input when needed. Let $r^{a}(\theta)=\frac{\pi^{a}_{\theta}(a|x)}{\pi^{a}_{\theta_{\text{old}}}(a|x)}$ and $r^{b}(\theta)=\frac{\pi^{b}_{\theta}(b|x)}{\pi^{b}_{\theta_{\text{old}}}(b|x)}$, where $\theta_{\text{old}}$ is the policy parameter before the update, and define the clipped surrogate objective terms, $L^{a}(\theta)=\mathop{\mathbb{E}}[\min(r^{a}(\theta)\hat{A},\text{clip}(r^{a}(\theta),1-\epsilon,1+\epsilon)\hat{A})],L^{b}(\theta)=\mathop{\mathbb{E}}[\min(r^{b}(\theta)\hat{A}$, $\text{clip}(r^{b}(\theta),1-\epsilon,1+\epsilon)\hat{A})]$, where $\hat{A}$ is the estimated advantage function. We aim to maximize the combined objective $L^{a}(\theta)+L^{b}(\theta)$, and perform the policy update as PPO does. By doing so, our policy learns the neural network hyperplane that minimizes long term constraint violation while suppressing the conservativeness of the hyperplane.
@@ 5 Experiments
. The proposed methods in the paper are each referred to as Oracle DH (DH based on Theorem 2, solved with support vector machine for each state), SL-DH (the supervised learning approach in Section 4.1), and RL-DH (the RL approach in Section 4.2). Through these experiments, we want to highlight
@@
    compatibility of DH with various representations of the safe domain $S$, based on HJ reachability in Xue et al.(2023)Xue, Zhan, Fränzle, Wang, and Liu, Lyapunov function, and manual design,
@@
    usage of the SL-DH and RL-DH for safe RL of versatile tasks, and comparison against unconstrained PPO (Schulman et al.(2017)Schulman, Wolski, Dhariwal, Radford, and Klimov), and PPO-Lagrangian (Ray et al.(2019)Ray, Achiam, and Amodei).
@@
1. Moore-Greitzer jet engine (Oracle DH vs. SL-DH, Figure 1(a)): The dynamics of this system is provided in Xue et al.(2023)Xue, Zhan, Fränzle, Wang, and Liu ($n$=2, $m$=1), in which the maximal control invariant set of the target constraint set $X=\{x|\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.5\}$ is computed by the discounted infinite-horizon HJ reachability formulation. The noticeable feature of this value function is that it is flat and zero everywhere inside the verified invariant set $S$. Thus, the safety constraint cannot be obtained from the value function, where the DH can be of great use. $\pi_{\text{ref}}$ is designed to stabilize to the origin, which can exit $S$ and violate safety. We use the computed $S$ to train the SL-DH, and compare the closed-loop trajectories under (a) $\pi_{\text{ref}}$, and safety filters based on (b) the Oracle DH, and (c) the SL-DH. The SL-DH safety filter closely approximates the Oracle DH safety filter.
2. Inverted Pendulum (CBF vs. SL-DH, Figure 1(b)): The reference controller $\pi_{\text{ref}}$ bounces between extreme torques and turns into a stabilizing controller after 12s. It easily exits the target angle region $X\!=\!\{x|\;|\theta|\!<\!0.3\}$. A level set of a quadratic Lyapunov function in $X$ is chosen as $S$. A CBF $h$ is also derived from this Lyapunov function. The dynamics of this system and $\pi_{\text{ref}},S,h$ are detailed in Wabersich et al.(2023)Wabersich, Taylor, Choi, Sreenath, Tomlin, Ames, and Zeilinger. Both CBF and SL-DH filters successfully maintain the system within the safe set while intervening $\pi_{\text{ref}}$ smoothly before the trajectory hits the boundary.
3. Kinematic Vehicle (Effect of $\Delta t$ on SL-DH, Figure 1(c, d)): The state of the vehicle consists of x and y position, heading, and velocity, while the input consists of the yaw rate and acceleration ($n$=4, $m$=2). The reference controller is a feedback controller designed to navigate to a goal point. The control invariant set and the CBF is hand-designed to avoid entering an unsafe region centered at the origin. More details are available in Appendix. We train SL-QP with multiple lookahead times: $\Delta t=0.1,0.5,1$. While $\pi_{\text{ref}}$ passes through the unsafe region, safety filters based on SL-DH and CBF are able to prevent the vehicle from entering it. A lower lookahead time leads to a less restrictive but more myopic intervention of the filter, resulting in trajectories approaching closer to the unsafe region. With a larger lookahead time, the safety filter is more preemptive with respect to the boundary. We also note that a higher lookahead time can be beneficial for the learning in that the resulting DH varies more smoothly with respect to state. However, an excessively high value can result in the safety filter to be overly conservative.
@@
4. Cart-Pole (Safe RL for various tasks, Figure 2): We show the utility of SL-DH and RL-DH for safely learning various tasks, subjected to the identical target constraint. Specifically, we compare PPO, whose policy is filtered by SL-DH and RL-DH during its task training, to unconstrained PPO and PPO-Lagrangian. The CartPole environment is defined in Towers et al.(2023)Towers, Terry, Kwiatkowski, Balis, Cola, Deleu, Goulão, Kallinteris, KG, Krimmel, Perez-Vicente, Pierré, Schulhoff, Tai, Shen, and Younis ($n=4,m=1$). The considered target constraint is $X\!=\!\{x|\;|s|\leq 0.5\}$, where $s$ is the position of the cart. The classic task is to maximize the length of the trajectory before termination (Task 1). We define two new tasks: stabilizing the cart at $s=0.4$ (Task 2), and maximizing cart speed while avoiding termination (Task 3). These new tasks challenge to approach the constraint’s boundary more than the classic task. Training of the SL-DH and RL-DH is detailed in Appendix.
Both SL-DH and RL-DH filtered PPOs demonstrate competitive performance compared to unconstrained PPO. Moreover, SL-DH shows no constraint violations, while RL-DH exhibits significantly fewer violations than PPO-Lagrangian across all tasks. RL-DH also achieved zero constraint violations in many instances. We also discovered that our method results in lower constraint violations across a broader range of initial states in the state space, as visualized in Appendix. Although the PPO-Lagrangian eventually learns to address safety constraints and achieve the task, the training is slower and violates safety more compared to our methods. This instability is mainly caused by the simultaneous learning of performance and safety, associated with unstable updates of the Lagrange multiplier.
@@ 6 Conclusion and Future Work
In this work, we have proposed a novel learning-based approach for the design of safety filters, focusing primarily on the control input constraint, which is key to ensuring safety. Our method, which employs a neural network to produce parameters of a discriminating hyperplane, offers a more general viewpoint of how to construct safety filters for general control-affine systems. We have shown that our approach is modular, working together with any control invariant set representation, and also compatible with any performance-driven objectives, thus replacing the need for safe RL approaches that combines performance and safety. While we have shifted the focus from the traditional certificate functions, we recognize that they still play an important role in verifying control invariant sets. Looking ahead, enhancing the robustness of the RL-DH safety filter against potential distribution shifts will be crucial for the improvement of our RL-based approach. Furthermore, we are keen on extending of our method for uncertain systems, as discussed in Lopez et al.(2020)Lopez, Slotine, and How; Brunke et al.(2022b)Brunke, Zhou, and Schoellig; Cohen et al.(2023)Cohen, Mann, Leahy, and Belta, which may include extensions to non-affine forms of the control input constraints like second-order cone constraints, as in Castañeda et al.(2021)Castañeda, Choi, Zhang, Tomlin, and Sreenath; Dhiman et al.(2021)Dhiman, Khojasteh, Franceschetti, and Atanasov; Taylor et al.(2021)Taylor, Dorobantu, Dean, Recht, Yue, and Ames, and to multiple constraints, as in Kim et al.(2023a)Kim, Lee, and Oh.
@@ 7.1 Proof of Theorem 1
We prove Theorem 1 by proving Lemma 2, since from Definitions 1 and 3, Lipschitz continuity of $\pi_{\text{ref}}$ implies that $a(x)^{\top}u\geq b(x)$ is a valid safety constraint, and thus, $\{x|a(x)^{\top}u=b(x)\}$ being a discriminating hyperplane. Under the conditions in Theorem 1, (5) can be written as a QP problem which satisfies the conditions in Morris et al.(2013)Morris, Powell, and Ames, which states that its solution is unique and Lipschitz w.r.t. $x$. $\square$
@@ 7.3 Implementation Details
SL-DH: In the loss function, we typically set $\gamma_{\text{pos}}>\gamma_{\text{neg}}$, since in order for the learned hyperplane to be a valid discriminating hyperplane, it is imperative to not misclassify unsafe inputs. Next, due to prediction error of the neural-network output, the direct use of raw predictions for the safety constraint $a_{\theta}(x)^{\top}u\geq b_{\theta}(x)$ might not ensure safety. To improve safety, we employ a minor adjustment factor, $\epsilon\geq 0$, to tighten the predicted hyperplane constraint. Specifically, we use the revised constraint, $a_{\theta}(x)^{\top}u\geq b_{\theta}(x)+\epsilon$. This adjustment can be viewed as a straightforward calibration step, a practice commonly seen in many deep neural network applications Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger. It should be noted, however, that over-utilizing this calibration (i.e., setting an excessively high value for $\epsilon$) can lead to overly conservative behavior by the safety filter. Nevertheless, in our experimental results, we observe no such over-conservatism, indicating minimal, yet effective calibration.

### v2

@@ 3.3 Sample-and-hold lookahead-based discriminating hyperplane
For $\Delta t>0$, there exists $a:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and $b:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that are Lipschitz continuous in $x$, such that $\Pi(x)=\{u\in U\;|\;a(x)^{\top}u\geq b(x)\}\subseteq\{u\in U\;|\;I_{S}(\mathrm{x}(\Delta t))=1\}$ where $\mathrm{x}(\cdot)$ is from (1) with $\mathrm{x}(0)=x,\mathrm{u}(\cdot)\equiv u$. Moreover, for small enough $\Delta t$, $\Pi(x)$ is not empty if $\{u\in U\;|\;I_{S}(\mathrm{x}(\Delta t))=1\}$ has a non-empty interior. (Proof: Appendix 7.2)
By using the lookahead approach, $\Pi(x)$ satisfying Theorem 2 anticipates into the future, which is also the main mechanism of CBFs (Choi et al., 2023) or predictive filter (Wabersich and Zeilinger, 2021). If the lookahead time is small, then the constraint would be active only on states that are close to the boundary. In contrast, larger $\Delta t$ will encode additional safety margin from the boundary of $S$, however, it might make $\Pi(x)$ an empty set. In Section 5, we demonstrate how this design results in a “smooth braking” behavior as $\mathrm{x}(\cdot)$ approaches the boundary of the safe domain, similarly to the CBF-based design, and how different choice of the lookahead time affects this behavior.
4 Learning Discriminating Hyperplane
This section presents supervised and reinforcement learning approaches to learn the discriminating hyperplane from black-box system trajectory data.
4.1 Supervised learning approach
In our supervised learning approach, we use the lookahead-based discriminating hyperplane in Section 3.3 to construct a training label for the neural network that learns the hyperplane. We train $[a_{\theta}(x),b_{\theta}(x)]\in\mathbb{R}^{m+1}$, where $\theta$ is the neural network weights, in a supervised manner using a dataset of state and input transition pairs, and their labels detailing if an input is safe. Specifically, $N$ states, $\{x_{i}\}_{i=1}^{N}$, are uniformly sampled from the control invariant set $S$. Then $M$ inputs per state, sampled uniformly from $U$, $\{u_{ij}\}_{j=1}^{M}$, are applied to the dynamical system by sample-and-hold for the lookahead time $\Delta t$, resulting in the terminal state $\mathrm{x}_{ij}(\Delta t)$. The indicator function of the set $S$ at the terminal state, $I_{S}(\mathrm{x}_{ij}(\Delta t))$, determines the label of whether the next state is safe or unsafe:
@@ 4.1 Supervised learning approach
If the prediction of safety from the neural network matches the labels, the loss is not incurred. False positive samples whose unsafe control input is predicted to be safe are counted in the first term, and false negative samples whose safe control input is predicted to be unsafe are counted in the second term, where $\gamma_{\text{pos}}$ and $\gamma_{\text{neg}}$ are weights for false positive and negative samples. We typically set $\gamma_{\text{pos}}>\gamma_{\text{neg}}$ as it is crucial to rule out unsafe inputs correctly.
@@ 4.2 Reinforcement learning approach
An alternative approach is to employ RL, when it is hard to find a control invariant set $S$ for the supervised learning method. We propose a method analogous to PPO, where instead of learning a policy that maximizes reward, we use an actor parametrizing the discriminating hyperplane to minimize instances of safety violation. Our method uses the form $\pi^{a}_{\theta}(a|x),\pi^{b}_{\theta}(b|x)$ as two normal distributions to sample the hyperplane parameters. During the rollout, based on the sampled $(a,b)$ for each $x$, with probability $1\!-\!\delta$, we sample $u$ from $\{u\in U|\;a^{\top}u\geq b\}$ which enforces the learned safety constraint, and with probability $\delta$, we sample $u$ from $U$, allowing an exploration with a small probability. Given the target constraint set $X$, the reward for learning the hyperplane is designed as
$r(x,u)=\begin{cases}1+d(u)&\text{if }x^{\prime}\in X,\\ c&\text{otherwise,}\end{cases}$ (7)
where $x^{\prime}$ is the next state after taking action $u$, $c\leq 0$ induces a negative reward when safety is violated, as $x^{\prime}\not\in X$, and $d(u)\geq 0$ is an optional bonus term that can be added to the reward if the input $u$ is in the action space $U$. The bonus term can be necessary when the action space is small, which leads to the actor’s predicted hyerplanes being outside of the action space. The reward is positive when safety is satisfied and the bonus term incentivizes the hyperplane to only constrain the input when needed. Let $r^{a}(\theta)=\frac{\pi^{a}_{\theta}(a|x)}{\pi^{a}_{\theta_{\text{old}}}(a|x)}$ and $r^{b}(\theta)=\frac{\pi^{b}_{\theta}(b|x)}{\pi^{b}_{\theta_{\text{old}}}(b|x)}$, where $\theta_{\text{old}}$ is the policy parameter before the update, and define the clipped surrogate objective terms, $L^{a}(\theta)=\mathop{\mathbb{E}}[\min(r^{a}(\theta)\hat{A},\text{clip}(r^{a}(\theta),1-\epsilon,1+\epsilon)\hat{A})],L^{b}(\theta)=\mathop{\mathbb{E}}[\min(r^{b}(\theta)\hat{A}$, $\text{clip}(r^{b}(\theta),1-\epsilon,1+\epsilon)\hat{A})]$, where $\hat{A}$ is the estimated advantage function. We aim to maximize the combined objective $L^{a}(\theta)+L^{b}(\theta)$, and perform the policy update as PPO does. By doing so, our policy learns the neural network hyperplane that minimizes long term constraint violation while suppressing the conservativeness of the hyperplane.
@@ 5 Experiments
. The proposed methods in the paper are each referred to as Oracle DH (DH based on Theorem 2 in Section 3.3, solved with support vector machine for each state), SL-DH (the supervised learning approach in Section 4.1), and RL-DH (the RL approach in Section 4.2). Through these experiments, we want to highlight
@@
    compatibility of DH with various representations of the safe domain $S$, based on HJ reachability in Xue et al. (2023), Lyapunov function, and manual design,
@@
    usage of the SL-DH and RL-DH for safe RL of versatile tasks, and comparison against unconstrained PPO (Schulman et al., 2017), and PPO-Lagrangian (Ray et al., 2019).
@@
1. Moore-Greitzer jet engine (Oracle DH vs. SL-DH, Figure 1(a)): The dynamics of this system is provided in Xue et al. (2023) ($n$=2, $m$=1), in which the maximal control invariant set of the target constraint set $X=\{x|\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.5\}$ is computed by the discounted infinite-horizon HJ reachability formulation. The noticeable feature of this value function is that it is flat and zero everywhere inside the verified invariant set $S$. Thus, the safety constraint cannot be obtained from the value function, where the DH can be of great use. $\pi_{\text{ref}}$ is designed to stabilize to the origin, which can exit $S$ and violate safety. We use the computed $S$ to train the SL-DH, and compare the closed-loop trajectories under (a) $\pi_{\text{ref}}$, and safety filters based on (b) the Oracle DH, and (c) the SL-DH. The SL-DH safety filter closely approximates the Oracle DH safety filter.
2. Inverted Pendulum (CBF vs. SL-DH, Figure 1(b)): The reference controller $\pi_{\text{ref}}$ bounces between extreme torques and turns into a stabilizing controller after 12s. It easily exits the target angle region $X\!=\!\{x|\;|\theta|\!<\!0.3\}$. A level set of a quadratic Lyapunov function in $X$ is chosen as $S$. A CBF $h$ is also derived from this Lyapunov function. The dynamics of this system and $\pi_{\text{ref}},S,h$ are detailed in Wabersich et al. (2023). Both CBF and SL-DH filters successfully maintain the system within the safe set while intervening $\pi_{\text{ref}}$ smoothly before the trajectory hits the boundary. This demonstrates how SL-DH is able to achieve similar behavior to CBF without utilization of known system dynamics, assuming access to the safe set beforehand.
3. Kinematic Vehicle (Effect of $\Delta t$ on SL-DH, Figure 1(c, d)): The state of the vehicle consists of x and y position, heading, and velocity, while the input consists of the yaw rate and acceleration ($n$=4, $m$=2). The reference controller is a feedback controller designed to navigate to a goal point. The control invariant set and the CBF is hand-designed to avoid entering an unsafe region centered at the origin. More details are available in Appendix. We train SL-DH with multiple lookahead times: $\Delta t=0.1,0.5,1$. While $\pi_{\text{ref}}$ passes through the unsafe region, safety filters based on SL-DH and CBF are able to prevent the vehicle from entering it. A lower lookahead time leads to a less restrictive but more myopic intervention of the filter, resulting in trajectories approaching closer to the unsafe region. With a larger lookahead time, the safety filter is more preemptive with respect to the boundary. We also note that a higher lookahead time can be beneficial for the learning in that the resulting DH varies more smoothly with respect to state. However, an excessively high value can result in the safety filter to be overly conservative.
@@
4. Cart-Pole (Safe RL for various tasks, Figure 2): We show the utility of SL-DH and RL-DH for learning various tasks, subjected to the identical target constraint. Specifically, we compare PPO, whose policy is filtered by SL-DH and RL-DH during its task training, to unconstrained PPO and PPO-Lagrangian. The CartPole environment is defined in Towers et al. (2023) ($n=4,m=1$). The considered target constraint is $X\!=\!\{x|\;|s|\leq 0.5\}$, where $s$ is the position of the cart. The classic task is to maximize the length of the trajectory before termination (Task 1). We define two new tasks: stabilizing the cart at $s=0.4$ (Task 2), and maximizing cart speed while avoiding termination (Task 3). These new tasks challenge to approach the constraint’s boundary more than the classic task. Training of the SL-DH and RL-DH is detailed in Appendix. Both SL-DH and RL-DH filtered PPOs demonstrate competitive performance compared to unconstrained PPO. Moreover, SL-DH shows no constraint violations, while RL-DH exhibits significantly fewer violations than PPO-Lagrangian across all tasks. RL-DH also achieved zero constraint violations in many instances. We also discovered that our method results in lower constraint violations across a broader range of initial states in the state space, as visualized in Appendix. The training of the PPO-Lagrangian is slower and violates safety more compared to our methods, mainly caused by the simultaneous learning of performance and safety, associated with unstable updates of the Lagrange multiplier.
@@ 6 Conclusion and Future Work
In this work, we have proposed a novel learning-based approach for the design of safety filters, focusing primarily on the control input constraint, which is key to ensuring safety. Our method, which employs a neural network to produce parameters of a discriminating hyperplane, offers a more general viewpoint of how to construct safety filters for general control-affine systems. We have shown that our approach is modular, working together with any control invariant set representation, and also compatible with any performance-driven objectives, thus replacing the need for safe RL approaches that combines performance and safety. While we have shifted the focus from the traditional certificate functions, we recognize that they still play an important role in verifying control invariant sets. Looking ahead, enhancing the robustness of the RL-DH safety filter against potential distribution shifts will be crucial for the improvement of our RL-based approach. Furthermore, we are keen on extending of our method for uncertain systems, as discussed in Lopez et al. (2020); Brunke et al. (2022b); Cohen et al. (2023), which may include extensions to non-affine forms of the control input constraints like second-order cone constraints, as in Castañeda et al. (2021); Dhiman et al. (2021); Taylor et al. (2021), and to multiple constraints, as in Kim et al. (2023a).
@@ 7.1 Proof of Theorem 1
We prove Theorem 1 by proving Lemma 2, since from Definitions 1 and 3, Lipschitz continuity of $\pi_{\text{ref}}$ implies that $a(x)^{\top}u\geq b(x)$ is a valid safety constraint, and thus, $\{x|a(x)^{\top}u=b(x)\}$ being a discriminating hyperplane. Under the conditions in Theorem 1, (5) can be written as a QP problem which satisfies the conditions in Morris et al. (2013, Thm.1), which states that its solution is unique and Lipschitz w.r.t. $x$. $\square$
@@ 7.3 Implementation Details
Oracle-DH: For the Oracle-DH method, we employ Support Vector Machines (SVMs) to determine a discriminating hyperplane at every state along the trajectory. Inputs labeled as ‘safe’ or ‘unsafe’ based on the sample-and-hold approach described in Section 3.3 are used to train the SVM. The SVM output is then directly utilized as the discriminating hyperplane in the safety filter.
SL-DH: In the loss function, we typically set $\gamma_{\text{pos}}>\gamma_{\text{neg}}$, since in order for the learned hyperplane to be a valid discriminating hyperplane, it is imperative to not misclassify unsafe inputs. Next, due to prediction error of the neural-network output, the direct use of raw predictions for the safety constraint $a_{\theta}(x)^{\top}u\geq b_{\theta}(x)$ might not ensure safety. To improve safety, we employ a minor adjustment factor, $\epsilon\geq 0$, to tighten the predicted hyperplane constraint. Specifically, we use the revised constraint, $a_{\theta}(x)^{\top}u\geq b_{\theta}(x)+\epsilon$. This adjustment can be viewed as a straightforward calibration step, a practice commonly seen in many deep neural network applications Guo et al. (2017). It should be noted, however, that over-utilizing this calibration (i.e., setting an excessively high value for $\epsilon$) can lead to overly conservative behavior by the safety filter. Nevertheless, in our experimental results, we observe no such over-conservatism, indicating minimal, yet effective calibration.